### v1

# Bubble in bubbles

Abstract
We investigate the evolution and formation of double-layered vacuum bubbles during cosmological phase transitions with multiple vacua. Using a semiclassical approach with initial velocity fluctuations, we demonstrate that under certain conditions, quantum effects do not lead to the formation of double-layered vacuum bubbles, while flyover transitions allow for their stable formation by overcoming successive potential barriers. The evolution of these bubbles, including wall acceleration, collisions, and the formation of trapped regions, is explored through numerical simulations. Our results show that the dynamics of double-layered bubbles differ significantly from standard single-wall bubbles, with implications for cosmological observables such as gravitational wave production and baryogenesis. These findings indicate that flyover transitions represent a complementary decay mechanism to the conventional quantum tunneling process.
I Introduction
Phase transitions in the early Universe play a crucial role in shaping its subsequent evolution and the formation of cosmological structures [1, 2]. Traditionally, such transitions have been studied in the context of quantum tunneling, where a scalar field initially trapped in a metastable vacuum can tunnel through a potential barrier to reach a lower-energy vacuum [3, 4]. This process provides a quantum-mechanical for bubble nucleation, and its consequences have been widely investigated in both flat and curved spacetimes [5]. Quantum tunneling not only determines the decay rate of metastable vacua, but also underlies a range of cosmological phenomena, including first-order phase transitions, baryogenesis [6, 7, 8], and the generation of stochastic gravitational wave backgrounds [9, 10, 11, 12, 13, 14].
Beyond pure quantum tunneling, both semiclassical and classical mechanisms can also drive vacuum decay  [15, 16, 17, 18, 19, 20, 21]. Semiclassical mechanisms, sometimes referred to as flyover vacuum decay [16], involve the scalar field overcoming potential barriers due to quantum or thermal fluctuations, rather than tunneling through them. Classical vacuum decay occurs in multi-vacuum scenarios, where it is typically triggered by collisions between vacuum bubbles formed through quantum tunneling. When these bubbles collide, they may generate enough energy to push the scalar field over potential barriers, leading to the formation of new bubbles in deeper vacua, without the need for additional quantum tunneling. Both semiclassical and classical decay mechanisms have been studied in various contexts, including stochastic methods and numerical simulations in one-dimensional models, providing insights into the decay rates, bubble dynamics, and the role of non-quantum fluctuations.
In models with multiple metastable vacua, the landscape becomes richer, allowing for the coexistence of quantum tunneling, flyover, and classical vacuum decay mechanisms. Depending on the potential structure and the physical conditions of the early Universe, a metastable vacuum can decay via one or more of these processes, leading to complex bubble configurations, such as double-layered vacuum bubbles. These multi-step or multi-channel decay processes can significantly modify the dynamics of phase transitions, influencing wall-related physics, collisions, and the resulting field configurations.
In this work, we investigate flyover vacuum decay in scalar field theories with multiple vacua, focusing on the formation, evolution, and interactions of double-layered vacuum bubbles. We adopt a semiclassical approach with initial velocity fluctuations, and numerically simulate bubble formation, collisions, and the formation of trapped regions. Our study aims to characterize the distinct dynamics of double-layered bubbles, compare them with standard single-wall transitions, and explore their implications for cosmological observables such as gravitational waves. The paper is organized as follows. In Sec. II, we introduce the mechanism for the formation of double-layered bubbles and the semiclassical approach. In Sec. III, we present numerical results on the formation and evolution of double-layered bubbles. Sec. IV provides the conclusion and discussion.
II Mechanism for the formation of double-layered bubbles
To investigate the double-layered vacuum bubbles formation, we consider the vacuum decay arising from semiclassical approach [15],we first choose a test potential with three vacua as shown in Fig. 1,
$V(\phi)=\frac{\lambda}{4}\phi^{2}\left(\phi^{2}-\phi_{0}^{2}\right)^{2}+\epsilon\phi_{0}^{5}\left(\phi-\phi_{0}\right)+\alpha\phi^{2}\phi_{0}^{2}.$ (1)
Figure 1: Potential Eq. (1) with three minima at $\phi_{1}$, $\phi_{2}$ and $\phi_{3}$. The heights of the two potential barriers are labeled as $V_{b1}$and $V_{b2}$, and the potential differences between the adjacent minima are denoted by $\Delta V_{12}$ and $\Delta V_{23}$.
Here, the parameter $\epsilon\phi_{0}^{2}$ must be sufficiently small to ensure the existence of potential barriers. When $\epsilon\phi_{0}^{2}=0$, the potential exhibits a $Z_{2}$ symmetry. The dimensionless parameter $\alpha$ governs the difference between $\Delta V_{12}$ and $\Delta V_{23}$, typically $\Delta V_{12}\approx\Delta V_{23}$ at $\alpha=0$ .
If the scalar field $\phi$ starts off localized in a metastable minimum at $\phi=\phi_{1}$, the surrounding potential barrier prevents classical evolution toward lower-energy vacua, yet still permits quantum tunneling to these vacua. In real space, such tunneling events correspond to the nucleation of vacuum bubbles whose interiors reside in the $\phi=\phi_{2}$ or $\phi=\phi_{3}$ vacuum. In the potential under consideration, quantum tunneling can give rise to three distinct types of vacuum bubbles: $\phi_{1}\to\phi_{2},\phi_{2}\to\phi_{3}$, and $\phi_{1}\to\phi_{3}$. The tunneling rate is approximately $\Gamma\approx e^{-S_{\mathrm{E}}}$, where $S_{\mathrm{E}}$ is the Euclidean action[22, 23, 24],
$S_{\mathrm{E}}=\int\mathrm{d}^{4}x\left[\frac{1}{2}(\partial_{\tau}\phi)^{2}+\frac{1}{2}(\nabla\phi)^{2}+V(\phi)\right].$ (2)
where $\tau$ is the Euclidean time. The dominant contribution comes from the path that minimizes $S_{\mathrm{E}}$, corresponding to the O(4)-symmetric bounce solution of
$\partial_{r_{\mathrm{E}}}^{2}\phi+\frac{3}{r_{\mathrm{E}}}\partial_{r_{\mathrm{E}}}\phi=V^{\prime}(\phi),$ (3)
with $r_{\mathrm{E}}^{2}=\tau^{2}+x^{2}+y^{2}+z^{2}$ and the boundary conditions $\partial_{r_{\mathrm{E}}}\phi=0$ at $r_{\mathrm{E}}=0$ and $\phi\to 0$ as $r_{\mathrm{E}}\to\infty$.
Figure 2: Representative vacuum bubble configuration arising through a quantum tunneling process. The potential parameters are chosen as $\alpha=0.05$, $\phi_{0}=1.5$, and $\epsilon\phi_{0}^{2}=0.068$. Under these conditions, the scalar field tunnels from the false vacuum into a true vacuum region, forming a bubble with a characteristic profile determined by the bounce solution.
In Fig. 2, we use FindBounce [25] to numerically obtain the critical bubble configurations $\phi_{c}$ for the three tunneling channels. For the $\phi_{1}\to\phi_{2}$ and $\phi_{2}\to\phi_{3}$ cases, the critical radius $R_{c}$ is defined as the radial position in the numerical profile [26]
$\phi_{c}\left(R_{\mathrm{c}}\right)=\frac{\Delta\phi}{2},$ (4)
where $\Delta\phi$ denotes the difference in field value separating the two local minima. It is important to note that the critical radius $R_{c}$ defines the threshold size beyond which a nucleated bubble will grow rather than collapse. It is interesting to note that the tunneling rate for $\phi_{1}\to\phi_{3}$ is comparable to that for $\phi_{1}\to\phi_{2}$, owing to the fact that $\Delta V_{12}>\Delta V_{23}$ [27]. In Fig. 2, the red dashed curve represents the critical bubble profile for the $\phi_{1}\to\phi_{3}$ transition. This profile can be naturally decomposed into two segments: one corresponding to the $\phi_{2}\to\phi_{3}$ region and the other to the $\phi_{1}\to\phi_{2}$ region,
$\phi_{c}=\begin{cases}\phi_{\mathrm{part1}},&\phi>\phi_{2},\\ \phi_{\mathrm{part2}},&\phi\leq\phi_{2}.\end{cases}$ (5)
To determine whether different segments of the $\phi_{1}\to\phi_{3}$ critical configuration can grow after nucleation, we compare their characteristic radii with the corresponding critical values $R_{c}^{12}$ and $R_{c}^{23}$. It can be seen that the critical radius of $\phi_{\mathrm{part2}}$ is equal to $R_{c}^{12}$. As a result, this part of the field configuration can successfully grow once nucleated. In contrast, the critical radius of $\phi_{\mathrm{part1}}$ is much smaller than $R_{c}^{23}$, causing this segment of the bubble to collapse. Consequently, the $\phi_{1}\to\phi_{3}$ tunneling process cannot be dynamically distinguished from the $\phi_{1}\to\phi_{2}$ transition. In practice, quantum tunneling nucleation effectively produces only $\phi_{1}\to\phi_{2}$ vacuum bubbles in this setup.
In addition to vacuum bubble nucleation via quantum tunneling, semiclassical mechanisms can also enable the scalar field to overcome the potential barrier and facilitate the phase transition. This process, known as flyover vacuum decay, essentially involves describing vacuum decay through a classical stochastic framework, where quantum or thermal fluctuations are modeled as initial fluctuations following a Gaussian distribution for a single real scalar field in 1+1 dimensions. Unlike quantum tunneling, in this scenario the scalar field acquires a non-zero configuration that trigger classically over the barrier, allowing bubble nucleation via a semiclassical pathway.
To study the subsequent evolution of the scalar field under this initial condition, we solve its equation of motion in the relevant cosmological setting. We consider phase transitions that occur on timescales much shorter than the Hubble time $H_{*}^{-1}$ at the epoch of the transition. In this regime, the effect of cosmic expansion is negligible, and the scalar field satisfies
$\square\phi-\frac{\mathrm{d}V}{\mathrm{d}\phi}=0,$ (6)
with $V$ being the scalar potential.
Regardless of the mechanism responsible for the formation of vacuum bubbles, we assume that they possess full spherical symmetry, such that the scalar field $\phi$ depends only on the radial coordinate $r$ and time $t$. Under this assumption, the standard Klein–Gordon equation in flat Minkowski spacetime takes the form
$\ddot{\phi}=\phi^{{\prime}{\prime}}+\frac{2}{r}\phi^{\prime}-\frac{\mathrm{d}V}{\mathrm{d}\phi},$ (7)
where dots and primes denote derivatives with respect to time $t$ and radius $r=\sqrt{x^{2}+y^{2}+z^{2}}$, respectively. This assumption enables a significant reduction in computational cost, making numerical simulations of vacuum bubble dynamics more tractable.
Unlike approaches that model flyover vacuum decay using initial field fluctuations, we follow the prescription of Ref. [16, 18], which instead introduces fluctuations in the velocity of the field while keeping the field itself at the local metastable vacuum value. This choice is motivated by the fact that representing the fluctuations as those of a free quantum field is more robust in this case, while still yielding results consistent with the field-fluctuation approach. Specifically, the scalar field is initially homogeneous at $\phi=\phi_{1}$, while its velocity profile occasionally acquires, through vacuum fluctuations, a localized Gaussian form
$\dot{\phi}(t=0,r)=A\exp\left(-\frac{r^{2}}{2R^{2}}\right).$ (8)
For such a fluctuation to seed an expanding bubble, its width must exceed the critical expansion radius can be estimated as $R_{c}\approx R_{0}=2\sigma/\Delta V_{12}$, where the wall tension is
$\sigma=\int_{\phi_{1}}^{\phi_{2}}\sqrt{2\left[V(\phi)-V{(\phi_{2})}\right]}\,\mathrm{d}\phi.$ (9)
Additionally, the amplitude $A$ must be at least $\sqrt{2V_{b1}}$ or larger in order to overcome the first potential barrier $V_{b1}$ via classical evolution. To uniformly specify, the value of $A$ at $\dot{\phi}(t=0,r=R)$ is defined as $A_{0}$ where it just equals $\sqrt{2V_{b1}}$.
This initial condition ansatz was originally proposed by Jose J. Blanco-Pillado et al. [16], and recent studies [28] have shown its consistency with numerical simulations of vacuum decay in $1+1$ dimensions starting from a metastable state. These simulations demonstrate that an initial Bose–Einstein distribution of fluctuations leads to bubble nucleation with Gaussian-distributed center-of-mass velocities, typically preceded by the formation of oscillons. Motivated by this agreement, we adopt the above initial condition in our analysis.
In our simulations, the $1+1$ dimensional setup uses a grid size of $12,000$, with a lattice length of $L/R_{0}=10$. In the $2+1$ dimensional setup, the grid consists of 500 points in each dimension, with $L/R_{0}=20$. In our simulations, we use a second-order finite difference scheme to discretize the equations. The time evolution is advanced using the leapfrog algorithm. We choose the timestep $\Delta t$ and lattice spacing $\Delta x$ such that $\Delta t=0.1\Delta x$, ensuring numerical stability and accuracy. This choice has been demonstrated in Refs. [26, 29] to provide excellent energy conservation properties.
Figure 3: Numerical simulation results with parameter values $A/A_{0}=1,R/R_{0}=0.8,\alpha=0$ (upper) and $A/A_{0}=1.2,R/R_{0}=1.1,\alpha=0$ (lower), respectively. The comparison illustrates how varying AA and $R$ affects the evolution of the scalar field configuration.
We consider two representative choices of the initial condition parameter with the corresponding evolution shown in Fig. 3. Vacuum transitions typically begin with the nucleation of $\phi_{2}$ bubbles within the $\phi_{1}$ background. For $A/A_{0}=1.2,R/R_{0}=1.1$, the subsequent evolution toward the true vacuum $\phi_{3}$—either via $\phi_{2}\to\phi_{3}$ or direct $\phi_{1}\to\phi_{3}$ transitions—depends sensitively on the relative vacuum energy gaps $\Delta V_{13}$ and $\Delta V_{23}$. If the $\phi_{2}$ vacuum is sufficiently long-lived, a two-step tunneling sequence may dominate the dynamics, driving the scalar field from the intermediate metastable state into the deeper true vacuum even when the tunneling rate to $\phi_{3}$ is highly suppressed. In the presence of three vacua, the flyover mechanism exhibits qualitatively distinct behavior compared to the standard two-vacuum scenario.
When the initial velocity is small enough, the dynamics of the scalar field are essentially indistinguishable between the two-vacuum and three-vacuum potentials, as the field cannot overcome either potential barrier. However, for sufficiently large initial velocities—particularly those with a Gaussian spatial profile—the central region of the perturbation, where the velocity is highest, can overshoot the intermediate vacuum $\phi_{2}$ and reach the true vacuum $\phi_{3}$, while the outer regions, having lower velocities, only transition to $\phi_{2}$. This results in the formation of a special double-layered vacuum bubbles structure, in which an inner bubble of $\phi_{3}$ is nested within an outer shell of $\phi_{2}$.
In this scenario, double-layered vacuum bubbles are generated via flyover transitions. The occurrence of such configurations depends on the dimensionless parameters $A/A_{0}$ and $R/R_{0}$, whose values are determined by model-dependent physical quantities such as the temperature and the scalar field mass. Within the framework of our toy model, we treat $A/A_{0}$ and $R/R_{0}$ as free parameters and perform a systematic scan over their values, as shown in Fig. 4. The nonlinear nature of the equations of motion leads to intricate structures—for instance, the appearance of a small orange region embedded within the green domain and a small green region within the orange area. Larger $A/A_{0}$ and $R/R_{0}$ correspond to greater initial kinetic energy, making it easier for the scalar field to overcome both potential barriers, whereas smaller values suppress such transitions. The resulting phase boundaries align with physical intuition and suggest the existence of critical values in both the $A/A_{0}$ and $R/R_{0}$ directions, beyond which vacuum decay can occur.
Figure 4: Entire parameter space where the blue region indicates the absence of a phase transition, the orange region suggests a transition expanding toward the $\phi_{2}$ vacuum, and the green region indicates a transition to the $\phi_{3}$ vacuum.
III Dynamic of double-layered bubbles
Having identified the parameter space for the formation of double-layered vacuum bubbles, we next investigate their stability. The stability of such configurations depends not only on the relative expansion velocities of the inner and outer walls—where a faster inner wall can overtake the outer wall, collapsing the double-layer into a single wall—but also on their behavior during collisions with other bubbles. Both effects can substantially influence the persistence and phenomenology of double-layered bubbles in the early Universe.
In the potential 1, increasing the parameter $\alpha$ lowers the values of $\Delta V_{23}$, while enhancing the potential difference between $\Delta V_{12}$. Since these changes can significantly affect the post-nucleation dynamics, tracking the subsequent bubble growth requires solving the full equation of motion derived from the Lagrangian [30].
$\ddot{r}+2\frac{1-\dot{r}^{2}}{r}=\frac{p}{\sigma}\left(1-\dot{r}^{2}\right)^{\frac{3}{2}},$ (10)
where $p=\Delta V_{12}~\text{or}~\Delta V_{23}$ denotes the pressure difference across the bubble wall.
To better understand the relationship between wall velocity and bubble radius, we simplify the equation of motion for the bubble wall. The Lorentz factor $\gamma$ characterizes the velocity of the bubble wall in relativistic terms, where $\gamma=1/\sqrt{1-v^{2}/c^{2}}$, and it accounts for the relativistic effects as the bubble expands. For an expanding bubble, the initial radius must exceed its critical radius. This condition can be solved analytically, with the initial condition $\gamma(R_{0})=1$, as derived in Ref. [30]. The equation is given by
$\gamma=\frac{pr}{3\sigma}+\frac{R_{0}^{2}}{r^{2}}-\frac{pR_{0}^{3}}{3\sigma r^{2}}\approx\frac{2r}{3R_{0}}+\frac{R_{0}^{2}}{3r^{2}},$ (11)
where in the last step, we assume that the initial radius is only slightly larger than the critical one, i.e., $R_{0}\approx R_{c}$. As shown in Fig. 5, the parameter $\alpha$ directly affects the potential differences $\Delta V_{12}$ and $\Delta V_{23}$, leading to distinct evolutionary outcomes. This can be understood from Eq. (11): the acceleration of a bubble wall is determined by the pressure difference across it. A larger potential difference drives the outermost wall to move faster, while the inner wall propagates more slowly. In contrast, a negative $\alpha$ can cause the inner wall to overtake the outer wall, resulting in a collision and subsequent merging of the two walls. Conversely, if the outer wall expands faster than the inner wall, the vacuum bubble can stably maintain a double-wall structure. Such a configuration may enhance or amplify wall-related physical phenomena, including gravitational wave production and baryogenesis.
Figure 5: Dynamics of vacuum decay with $\alpha=0.035$, $\phi_{0}=1.5$, $A/A_{0}=7$ and $R/R_{0}=0.59$ (top), and with $\alpha=-0.038$, $\phi_{0}=1.5$, $A/A_{0}=11$ and $R/R_{0}=0.2$ (lower), respectively.
Having discussed how the relative velocities of the inner and outer walls govern the stability of a single double-layered bubble, we now consider the scenario in which two such bubbles nucleate close enough to interact. Assuming that the bubbles are aligned along the $z$-axis, the system exhibits approximate cylindrical symmetry, which allows us to simplify the analysis of their subsequent evolution. In this setup, the dynamics of the scalar field during the collision can be described by the Klein–Gordon equation in cylindrical coordinates:
$\partial_{t}^{2}\phi-\partial_{r}^{2}\phi-\frac{1}{r}\partial_{r}\phi-\partial_{z}^{2}\phi=-\frac{\mathrm{d}V}{\mathrm{d}\phi},\quad r=\sqrt{x^{2}+y^{2}}.$ (12)
We place two identical Gaussian wave packets at positions $z/R_{0}=\pm D$, with $D$ chosen to be larger than the Gaussian width to prevent initial overlap or interference. As shown in Fig. 6, double-layered vacuum bubbles are formed and the collision occurs before any significant interaction between the walls. At $t/R_{0}\simeq 3.16$, the outer walls of the two bubbles begin to collide, which drives the regions near the bubble walls from the $\phi_{2}$ vacuum to the $\phi_{3}$ vacuum. Such elliptical vacuum bubbles formed through classical dynamics and the analysis of their formation have been discussed in [20, 21]. The key difference between the work presented here and their studies is that the colliding vacuum bubbles in this work originate from flyover vacuum decay.
The walls of these classically formed elliptical bubbles subsequently collide with the inner walls of the original bubbles, leading to the formation of so-called trapping regions [31, 32], where trapping at the false vacuum occurs after the collisions. The formation of these trapping regions depends sensitively on the wall thickness and typically occurs in collisions involving thin-walled bubbles. Later, interactions between trapping regions and inner regions at $t/D\simeq 4.73$ further convert trapped $\phi_{2}$ regions into the $\phi_{3}$ vacuum. As time progresses, all vacua will eventually settle into the $\phi_{3}$ state by the scalar radiation. From our numerical results, we find that the energy–momentum tensor associated with the dynamics of double-layered bubbles exhibits significant differences from that of standard single-wall bubble collisions. Consequently, these differences are expected to lead to distinct features in the power spectrum of gravitational wave energy density.
Figure 6: Double-layered vacuum bubbles are formed and the collision occurs before any significant interaction between the walls. The top row corresponds to $t/R_{0}=1.56$ and $t/R_{0}=3.16$, and the bottom row corresponds to $t/R_{0}=3.95$ and $t/R_{0}=4.73$, from left to right. The simulation parameters are $A/A_{0}=1.7$, $R/R_{0}=0.98$, $D/R_{0}=3.95$, and $\alpha=0$.
IV Conclusion and discussion
In this work, we have investigated flyover vacuum decay in a multi-vacuum scenario. We find that, for certain choices of initial parameters, this process can lead to the formation of double-layered vacuum bubbles. In contrast, double-layered bubbles generated purely by quantum tunneling are generally unstable and quickly decay into single-wall vacuum bubbles.
Our study employs a semiclassical approach originally introduced by Braden et al. [15] , who used the stochastic method to study vacuum decay numerically in 1+1 dimensions and demonstrated quantitative agreement with standard tunneling calculations. This raises the question of whether this semiclassical approach merely provides an alternative approximation for the decay rate or actually describes a distinct channel of vacuum decay. From our analysis of bubble nucleation in multi-vacuum potentials, the existence of stable double-layered bubbles suggests that flyover transitions indeed provide an independent decay channel, different from conventional quantum tunneling.
Furthermore, the dynamics of double-layered bubbles differ significantly from single-wall bubbles. The expansion and collisions of these bubbles modify the scalar field evolution, which in turn can drive the evolution of coupled fluids, potentially leading to distinct cosmological signatures such as gravitational waves or baryogenesis. Additionally, the presence of two types of nucleated bubbles introduces multiple possible decay pathways, giving rise to gravitational wave power spectra that depend on a variety of parameters controlling the bubble dynamics. Overall, our results indicate that semiclassical flyover transitions enrich the phenomenology of vacuum decay in multi-vacuum potentials, opening new possibilities for observational consequences in the early Universe.
acknowledgements
This work is supported in part by the National Key Research and Development Program of China No. 2020YFC2201501, in part by the National Natural Science Foundation of China No. 12475067 and No. 12235019.